How many terms of progression 3, 5, 7, 9 . . . must be taken

in order that their sum will be 2600?


1
Expert's answer
2021-09-15T02:58:33-0400

Let us use the following formula for the sum of "n" terms of arithmetic progression: "S_n=\frac{2a_1+d(n-1)}2n." In our case, "a_1=3,\ d=2,\ S_n=2600." It follows that "2600=\frac{6+2(n-1)}2n=(3+n-1)n=(n+2)n."

Taking into account that "2600=52\cdot 50," we conclude that "n=50." Therefore, 50 terms must be taken.